# Deciding Local Unitary Equivalence of Graph States in Quasi-Polynomial Time

### v1

@@ Acknowledgements
Recall that $G$ and $G^{\prime}$ are said LC-equivalent if there exist a sequence of (possibly repeating) vertices $a_{1},\cdots,a_{m}$ such that $G^{\prime}=G\star a_{1}\star\cdots\star a_{m}$. According to Bouchet, $G$ and $G^{\prime}$ are LC-equivalent if and only if there exist $A,B,C,D\subseteq V$ such that
@@
    $A\cap D\leavevmode\nobreak\ \Delta\leavevmode\nobreak\ B\cap C=V$
@@
Furthermore, $A\cap D\leavevmode\nobreak\ \Delta\leavevmode\nobreak\ B\cap C=V$.
Now, suppose that $G$ and $G^{\prime}$ are equivalent and there exists $A,B,C,D\subseteq V$ satisfying (i) and (ii). Applying a local complementation on some vertex $w$ in $G$ results in the graph $G\star w$, which is also LC-equivalent to $G^{\prime}$. Define
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|(D\Delta(N_{G}(w)\cap B))\cap\{u\}\cap N_{G^{\prime}}(v)|+|(C\Delta(N_{G}(w)\cap A))\cap\{u\}\cap\{v\}|$
@@
+|$\leavevmode\nobreak\ \leavevmode\nobreak\ +|\{w\}\cap D\cap N_{G}(u)\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap N_{G}(u)\cap\{v\}|$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|N_{G}(w)\cap B\cap\{u\}\cap N_{G^{\prime}}(v)|+|N_{G}(w)\cap A\cap\{u\}\cap\{v\}|\bmod 2$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|(A\Delta(\{w\}\cap C))\cap(N_{G}(u)\Delta N_{G}(w)\Delta\{u\})\cap\{v\}|$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|(D\Delta(N_{G}(w)\cap B))\cap\{u\}\cap N_{G^{\prime}}(v)|+|(C\Delta(N_{G}(w)\cap A))\cap\{u\}\cap\{v\}|$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|B\cap N_{G}(w)\cap N_{G^{\prime}}(v)|+|A\cap N_{G}(w)\cap\{v\}|+|D\cap\{w\}\cap N_{G^{\prime}}(v)|+|C\cap\{w\}\cap\{v\}|$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +|B\cap\{u\}\cap N_{G^{\prime}}(v)|+|A\cap\{u\}\cap\{v\}|+|B\cap\{u\}\cap N_{G^{\prime}}(v)|+|A\cap\{u\}\cap\{v\}|$
+|$\leavevmode\nobreak\ \leavevmode\nobreak\ +|\{w\}\cap D\cap N_{G}(w)\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap N_{G}(w)\cap\{v\}|$
+|$\leavevmode\nobreak\ \leavevmode\nobreak\ +|\{w\}\cap D\cap\{u\}\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap\{u\}\cap\{v\}|$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak$ $G\star w$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $G\star w$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak$ $G\wedge w_{1}w_{2}$
@@
  • $A^{\prime}=(A\cap\setminus\{w_{1},w_{2}\})\cup(C\cap\{w_{1},w_{2}\})$
  • $B^{\prime}=(B\cap\setminus\{w_{1},w_{2}\})\cup(D\cap\{w_{1},w_{2}\})$
  • $C^{\prime}=(C\cap\setminus\{w_{1},w_{2}\})\cup(A\cap\{w_{1},w_{2}\})$
  • $D^{\prime}=(D\cap\setminus\{w_{1},w_{2}\})\cup(B\cap\{w_{1},w_{2}\})$
Let us design an algorithm that produces a sequence of (possibly repeating) vertices $s=(a_{1},\cdots,a_{m})$ such that $G^{\prime}=G\star a_{1}\star\cdots\star a_{m}$. Initialise $G_{0}=G$, $s_{0}=[\leavevmode\nobreak\ ]$ an empty sequence of vertices and $A_{0}=A$, $B_{0}=B$, $C_{0}=C$, $D_{0}=D$.
@@
    If there is a vertex $u$ in case 2 or 6: let $s_{0}\leftarrow s_{0}+[u]$, $G_{0}\leftarrow G_{0}\star u$, $A_{0}\leftarrow A_{0}\Delta(\{w\}\cap C_{0})$, $B_{0}\leftarrow B_{0}\Delta(\{w\}\cap D_{0})$, $C_{0}\leftarrow C_{0}\Delta(N_{G_{0}}(w)\cap A_{0})$, $D_{0}\leftarrow D_{0}\Delta(N_{G_{0}}(w)\cap B_{0})$. Repeat until there is no vertex in case 2 or 6 left.
@@
    If there is a vertex $u$ in case 4 or 5: let $v\in N_{G_{0}}(u)$ such that $v$ is also in case 4 or 5. Let $s_{0}\leftarrow s_{0}+[u,v,u]$, $G_{0}\leftarrow G_{0}\wedge uv$, $A_{0}\leftarrow(A_{0}\cap\setminus\{u,v\})\cup(C_{0}\cap\{u,v\})$, $B_{0}\leftarrow(B_{0}\cap\setminus\{u,v\})\cup(D_{0}\cap\{u,v\})$, $C_{0}\leftarrow(C_{0}\cap\setminus\{u,v\})\cup(A_{0}\cap\{u,v\})$, $D_{0}\leftarrow(D_{0}\cap\setminus\{u,v\})\cup(B_{0}\cap\{u,v\})$. Then go to step 1.
@@
Termination. The number of vertices in case 1 or 3 stricly increases at each iteration of the algorithm.
@@ Appendix B Interpretation of the constraints in terms of local Clifford operators
$I$ if $v\in\overline{B}\cap\overline{C}$ $H$ if $v\in\overline{A}\cap\overline{D}$ $Z(\pi/2)$ if $v\in\overline{B}\cap C$ $Z(\pi/2)H$ if $v\in\overline{A}\cap{D}$ $X(\pi/2)$ if $v\in B\cap\overline{C}$ $X(\pi/2)H$ if $v\in{A}\cap\overline{D}$
@@
More details on the LC-equivalence of graphs and the corresponding Clifford operators can be found in [13, 18]. From a graph state point of view, Proposition 3.5 provides an efficient algorithm to decide whether two graph states are LC-equivalent under some constraints on the Clifford operators. Notice that such constraints should be expressible as a linear equation through the correspondence given in Table 1. We give below a non-exhaustive family of constraints expressible as linear equations (in the following, $k$ denotes an integer).
  • $v\notin B$$U_{v}$ is $Z(k\pi/2)$ up to Pauli;
  • $v\notin C$$U_{v}$ is $X(k\pi/2)$ up to Pauli;
  • $v\notin A$$U_{v}$ is $Z(k\pi/2)H$ up to Pauli;
  • $v\notin D$$U_{v}$ is $X(k\pi/2)H$ up to Pauli;
[…]
  • $v\in A$ iff $v\in D$$U_{v}$ is I, $X(\pi/2)$, $Z(\pi/2)$ or $H$ up to Pauli, i.e. $U_{v}^{2}$ is a Pauli;
  • $v\in A$ iff $v\in B$$U_{v}$ is $X(\pi/2)$ or $X(k\pi/2)H$ up to Pauli;
  • $v\in A$ iff $w\in A$, $v\in B$ iff $w\in B$, $v\in C$ iff $w\in C$, $v\in D$ iff $w\in D$$U_{v}=U_{w}$ up to Pauli.
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak$ $G\star u$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $G\star u$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ G\leavevmode\nobreak\ \leavevmode\nobreak$ $G\wedge uv$
@@
    If there is a vertex $u$ of type X in $G_{1}$ (resp. $G_{2}$) such that $\{u\}\cup N_{G_{1}}(u)$ (resp. $\{u\}\cup N_{G_{2}}(u)$) is not a minimal local set of dimension 1: find a minimal local set $M$ contained in $\{u\}\cup N_{G_{1}}(u)$ (resp. $\{u\}\cup N_{G_{2}}(u)$) and check that $M$ is a minimal local set of same dimension in both graph (if not, they are not LU-equivalent). If this is the case, add $M$ to $\mathcal{M}$ then go to step 5.
@@
Complexity. The time-complexity of the algorithm is given by the time-complexity of step 1, as it is asymptotically the most computationnally expensive step. The MLS cover can be computed in runtime $O(n^{6.38})$, where $n$ is the order of the graph. It should be noted that giving the type of a vertex with respect to some minimal local set can be done in time-complexity $O(n^{3})$. Indeed, in the case of a minimal local $L$ set of dimension 1, finding $D$ such that $L=D\cup Odd_{G}(D)$ reduces to finding the kernel of some matrix with coefficients in $\mathbb{F}_{2}$ (see details in [10]), which can be done using Gaussian elimination.
@@
In the main algorithm described in Section 3.3, the number of vertices of type $Z$ is directly linked to the runtime of the algorithm. Thus, it is preferable that the number of vertices of type Z is low. It should be noted that right after step 1, it is possible to reduce the number of vertices of type Z by adding a linear number of minimal local sets to $\mathcal{M}$. Indeed, following [11], if there exists a set $K$ of vertices of type Z of size more than $\lfloor n/2\rfloor+1$ where $n$ is the order of the graphs, then a minimal local set $L$ within $K$ can be found, and adding $L$ to $\mathcal{M}$ transforms at least one vertex of type $Z$ into a vertex of type $\bot$. Repeating the operations leads to graphs where the number of vertices of type $Z$ is at most $\lfloor n/2\rfloor$. Furthermore, notice than the remaining of the algorithm above may only decrease the number of vertices of type $Z$.
@@
Suppose $G_{1}$ and $G_{2}$ LCr-equivalent. According to Lemma 3.14, $G_{1}$ and $G_{2}$ are related by a single $r$-local complementation over a multiset $S$ whose support lies in $V_{X}$, along with a sequence of local complementations on the vertices of $V\setminus(V_{X}\cup V_{Z})$. We note $G_{2}=G_{1}\star^{r}S\star u_{1}\star\cdots u_{k}$. The edges toggled by an $r$-local complementation over $S$ are described by an element $\omega\in\Omega$. Let us decompose $\omega$ as a linear combination of basis vectors of $\Omega$: $\omega=\omega_{1}+\cdots+\omega_{t}$, where each $\omega_{i}\in\mathcal{B}$. In $G^{\#}_{1}$ and $G^{\#}_{2}$, let $V_{\omega}=\bigcup_{i\in[1,t]}\mathcal{P}^{\omega_{i}}$. Then, $G^{\#}_{2}=G^{\#}_{1}\star^{1}V_{\omega}\star u_{1}\star\cdots u_{k}$. Note that the 1-local complementation over the set $V_{\omega}$ corresponds to the composition of local complementations on each element of $V_{\omega}$. Thus, $G_{1}$ is mapped to $G_{2}$ by a sequence of local complementations that satisfy the additional constraints.
@@
According to Proposition 3.22, $S$ is genuine. Let $K_{\text{max}}$ be one of the biggest subset (by inclusion) of $V\setminus\textup{{supp}}(S)$ such that $\sum_{N_{G}(u)=K_{\text{max}}}S(u)$ is odd: then $S\bullet\Lambda_{G}^{K_{\text{max}}}$ is odd. If $|V\setminus\textup{{supp}}(S)|\leqslant r+1$, then $|K_{max}|\leqslant r+1$, contradicting the $r$-incidence of $S$. Thus $|V\setminus\textup{{supp}}(S)|\geqslant|K_{max}|\geqslant r+2$.
@@
$=\sum_{N_{G}(u)=K_{0}}S(u)+|\{K\subseteq V\setminus\textup{{supp}}(S)\leavevmode\nobreak\ |\leavevmode\nobreak\ K_{0}\subsetneq K\}|\bmod 2\text{\leavevmode\nobreak\ \leavevmode\nobreak\ by hypothesis of induction}$
@@
Let $S^{\prime}$ be the multiset obtained from $S$ by choosing, for every $K\subseteq V\setminus\textup{{supp}}(S)$ s.t. $|K|>1$, a single vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G}(u)=K$, and setting $S^{\prime}(u)=S(u)-1$ (the multiplicity of other vertices remain unchanged). Let us prove that $S^{\prime}$ is $r$-incident and $G\star^{r}S=G\star^{r}S^{\prime}$. First, $S^{\prime}$ is $r$-incident. Indeed, let an integer $k\in[0,r)$, let $K_{1}\subseteq V\setminus\textup{{supp}}(S^{\prime})$ be a set of size $k+2$, and let $k^{\prime}=k-|K_{1}\cap\textup{{supp}}(S)|$. $S\bullet\Lambda_{G}^{K_{1}}$ is a multiple of $2^{r-k^{\prime}-\delta(k^{\prime})}$ by $r$-incidence of $S$, so is $S^{\prime}\bullet\Lambda_{G}^{K_{1}}$, as $S^{\prime}\bullet\Lambda_{G}^{K_{1}}=S\bullet\Lambda_{G}^{K_{1}}-|\{K\subseteq V\setminus\textup{{supp}}(S)\leavevmode\nobreak\ |\leavevmode\nobreak\ K_{1}\setminus\textup{{supp}}(S)\subseteq K\}|=S\bullet\Lambda_{G}^{K_{1}}-2^{r-k^{\prime}}$. Then, if $u$ or $v$ is in $\textup{{supp}}(S)$, $u\sim_{G\star^{r}S}v\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ u\sim_{G\star^{r}S^{\prime}}v\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ u\sim_{G}v$. If $u,v\in V\setminus\textup{{supp}}(S)$:
$u\sim_{G\star^{r}S}v$ $\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \left(u\sim_{G}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ S\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \left(u\sim_{G}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ S^{\prime}\bullet\Lambda_{G}^{u,v}+2^{r}=2^{r-1}\bmod 2^{r}\right)$
$\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \left(u\sim_{G}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ S^{\prime}\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ u\sim_{G\star^{r}S^{\prime}}v$
Thus, $G\star^{r}S=G\star^{r}S^{\prime}$. Notice also that $S^{\prime}$ is not genuine. Thus, by Proposition 3.22 there exists an $S^{\prime\prime}$ such that $G\star^{r-1}S^{\prime\prime}=G\star^{r}S^{\prime}=G\star^{r}S$.

### v3

@@ Acknowledgements
Recall that $G$ and $G^{\prime}$ are said LC-equivalent if there exists a sequence of (possibly repeating) vertices $a_{1},\cdots,a_{m}$ such that $G^{\prime}=G\star a_{1}\star\cdots\star a_{m}$. According to Bouchet, $G$ and $G^{\prime}$ are LC-equivalent if and only if there exist $A,B,C,D\subseteq V$ such that
@@
    $A\cap D\penalty 10000\ \Delta\penalty 10000\ B\cap C=V$
@@
Furthermore, $A\cap D\penalty 10000\ \Delta\penalty 10000\ B\cap C=V$.
Now, suppose that $G$ and $G^{\prime}$ are equivalent and there exist $A,B,C,D\subseteq V$ satisfying (i) and (ii). Applying a local complementation on some vertex $w$ in $G$ results in the graph $G\star w$, which is also LC-equivalent to $G^{\prime}$. Define
@@
$\penalty 10000\ \penalty 10000\ +|(D\Delta(N_{G}(w)\cap B))\cap\{u\}\cap N_{G^{\prime}}(v)|+|(C\Delta(N_{G}(w)\cap A))\cap\{u\}\cap\{v\}|$
@@
$\penalty 10000\ \penalty 10000\ +|\{w\}\cap D\cap N_{G}(u)\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap N_{G}(u)\cap\{v\}|$
$\penalty 10000\ \penalty 10000\ +|N_{G}(w)\cap B\cap\{u\}\cap N_{G^{\prime}}(v)|+|N_{G}(w)\cap A\cap\{u\}\cap\{v\}|\bmod 2$
@@
$\penalty 10000\ \penalty 10000\ +|(A\Delta(\{w\}\cap C))\cap(N_{G}(u)\Delta N_{G}(w)\Delta\{u\})\cap\{v\}|$
$\penalty 10000\ \penalty 10000\ +|(D\Delta(N_{G}(w)\cap B))\cap\{u\}\cap N_{G^{\prime}}(v)|+|(C\Delta(N_{G}(w)\cap A))\cap\{u\}\cap\{v\}|$
@@
$\penalty 10000\ \penalty 10000\ +|B\cap N_{G}(w)\cap N_{G^{\prime}}(v)|+|A\cap N_{G}(w)\cap\{v\}|+|D\cap\{w\}\cap N_{G^{\prime}}(v)|+|C\cap\{w\}\cap\{v\}|$
$\penalty 10000\ \penalty 10000\ +|B\cap\{u\}\cap N_{G^{\prime}}(v)|+|A\cap\{u\}\cap\{v\}|+|B\cap\{u\}\cap N_{G^{\prime}}(v)|+|A\cap\{u\}\cap\{v\}|$
$\penalty 10000\ \penalty 10000\ +|\{w\}\cap D\cap N_{G}(w)\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap N_{G}(w)\cap\{v\}|$
$\penalty 10000\ \penalty 10000\ +|\{w\}\cap D\cap\{u\}\cap N_{G^{\prime}}(v)|+|\{w\}\cap C\cap\{u\}\cap\{v\}|$
@@
$\penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\$$G\star w$
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\ \penalty 10000\$$G\star w$
@@
$\penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\$$G\wedge w_{1}w_{2}$
@@
  • $A^{\prime}=(A\setminus\{w_{1},w_{2}\})\cup(C\cap\{w_{1},w_{2}\})$
  • $B^{\prime}=(B\setminus\{w_{1},w_{2}\})\cup(D\cap\{w_{1},w_{2}\})$
  • $C^{\prime}=(C\setminus\{w_{1},w_{2}\})\cup(A\cap\{w_{1},w_{2}\})$
  • $D^{\prime}=(D\setminus\{w_{1},w_{2}\})\cup(B\cap\{w_{1},w_{2}\})$
Let us design an algorithm that produces a sequence of (possibly repeating) vertices $s=(a_{1},\cdots,a_{m})$ such that $G^{\prime}=G\star a_{1}\star\cdots\star a_{m}$. Initialise $G_{0}=G$, $s_{0}=[\penalty 10000\ ]$ an empty sequence of vertices and $A_{0}=A$, $B_{0}=B$, $C_{0}=C$, $D_{0}=D$.
@@
    If there is a vertex $u$ in case 2 or 6: let $s_{0}\leftarrow s_{0}+[u]$, $G_{0}\leftarrow G_{0}\star u$, $A_{0}\leftarrow A_{0}\Delta(\{u\}\cap C_{0})$, $B_{0}\leftarrow B_{0}\Delta(\{u\}\cap D_{0})$, $C_{0}\leftarrow C_{0}\Delta(N_{G_{0}}(u)\cap A_{0})$, $D_{0}\leftarrow D_{0}\Delta(N_{G_{0}}(u)\cap B_{0})$. Repeat until there is no vertex in case 2 or 6 left.
@@
    If there is a vertex $u$ in case 4 or 5: let $v\in N_{G_{0}}(u)$ such that $v$ is also in case 4 or 5. Let $s_{0}\leftarrow s_{0}+[u,v,u]$, $G_{0}\leftarrow G_{0}\wedge uv$, $A_{0}\leftarrow(A_{0}\setminus\{u,v\})\cup(C_{0}\cap\{u,v\})$, $B_{0}\leftarrow(B_{0}\setminus\{u,v\})\cup(D_{0}\cap\{u,v\})$, $C_{0}\leftarrow(C_{0}\setminus\{u,v\})\cup(A_{0}\cap\{u,v\})$, $D_{0}\leftarrow(D_{0}\setminus\{u,v\})\cup(B_{0}\cap\{u,v\})$. Then go to step 1.
@@
Termination. The number of vertices in case 1 or 3 strictly increases at each iteration of the algorithm.
@@ Appendix B Interpretation of the constraints in terms of local Clifford operators
II if $v\in\overline{B}\cap\overline{C}$ $H$ if $v\in\overline{A}\cap\overline{D}$ $Z(\pi/2)$ if $v\in\overline{B}\cap C$ $Z(\pi/2)H$ if $v\in\overline{A}\cap{D}$ $X(\pi/2)$ if $v\in B\cap\overline{C}$ $X(\pi/2)H$ if $v\in{A}\cap\overline{D}$
@@
More details on the LC-equivalence of graphs and the corresponding Clifford operators can be found in [13, 18]. From a graph state point of view, Proposition˜3.5 provides an efficient algorithm to decide whether two graph states are LC-equivalent under some constraints on the Clifford operators. Notice that such constraints should be expressible as a linear equation through the correspondence given in Table˜1. We give below a non-exhaustive family of constraints expressible as linear equations (in the following, $k$ denotes an integer).
  • $v\notin B$: $U_{v}$ is $Z(k\pi/2)$ up to Pauli;
  • $v\notin C$: $U_{v}$ is $X(k\pi/2)$ up to Pauli;
  • $v\notin A$: $U_{v}$ is $Z(k\pi/2)H$ up to Pauli;
  • $v\notin D$: $U_{v}$ is $X(k\pi/2)H$ up to Pauli;
  • $v\notin\overline{B}\cap\overline{C}$: $U_{v}$ is a Pauli;
  • $v\notin\overline{A}\cap\overline{D}$: $U_{v}$ is $H$ up to a Pauli;
  • $v\in A$ iff $v\in D$: $U_{v}$ is I, $X(\pi/2)$, $Z(\pi/2)$ or $H$ up to Pauli, i.e. $U_{v}^{2}$ is a Pauli;
  • $v\in A$ iff $v\in B$: $U_{v}$ is $X(\pi/2)$ or $X(k\pi/2)H$ up to Pauli;
  • $v\in A$ iff $w\in A$, $v\in B$ iff $w\in B$, $v\in C$ iff $w\in C$, $v\in D$ iff $w\in D$: $U_{v}=U_{w}$ up to Pauli.
@@
$\penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\$$G\star u$
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\ \penalty 10000\$$G\star u$
@@
$\penalty 10000\ \penalty 10000\ G\penalty 10000\ \penalty 10000\$$G\wedge uv$
@@
    If there is a vertex $u$ of type X in $G_{1}$ (resp. $G_{2}$) such that $\{u\}\cup N_{G_{1}}(u)$ (resp. $\{u\}\cup N_{G_{2}}(u)$) is not a minimal local set of dimension 1: find a minimal local set $M$ contained in $\{u\}\cup N_{G_{1}}(u)$ (resp. $\{u\}\cup N_{G_{2}}(u)$) and check that $M$ is a minimal local set of same dimension in both graphs (if not, they are not LU-equivalent). If this is the case, add $M$ to $\mathcal{M}$ then go to step 5.
@@
Complexity. The time-complexity of the algorithm is given by the time-complexity of step 1, as it is asymptotically the most computationally expensive step. The MLS cover can be computed in runtime $O(n^{6.38})$, where $n$ is the order of the graph. It should be noted that giving the type of a vertex with respect to some minimal local set can be done in time-complexity $O(n^{3})$. Indeed, in the case of a minimal local $L$ set of dimension 1, finding $D$ such that $L=D\cup Odd_{G}(D)$ reduces to finding the kernel of some matrix with coefficients in $\mathbb{F}_{2}$ (see details in [10]), which can be done using Gaussian elimination.
@@
In the main algorithm described in Section˜3.3, the number of vertices of type $Z$ is directly linked to the runtime of the algorithm. Thus, it is preferable that the number of vertices of type Z is low. It should be noted that at any point of the algorithm, following [11], if there exists a set $K$ of vertices of type Z of size more than $\lfloor n/2\rfloor+1$, then a minimal local set $L$ within $K$ can be found, and adding $L$ to $\mathcal{M}$ transforms at least one vertex of type $Z$ into a vertex of type $\bot$. Repeating the operation leads to a graph where the number of vertices of type $Z$ is at most $\lfloor n/2\rfloor$. The algorithm can then be restarted from step 2. Notice that the number of vertices of type $\bot$ never decreases, thus this procedure will be performed less than $n$ times. Adding random minimal local sets to $\mathcal{M}$ is another way of trying to reduce the number of vertices of type Z.
@@
Suppose $G_{1}$ and $G_{2}$ LCr-equivalent. According to Lemma˜3.14, $G_{1}$ and $G_{2}$ are related by a single $r$-local complementation over a multiset $S$ whose support lies in $V_{X}$, along with a sequence of local complementations on the vertices of $V\setminus(V_{X}\cup V_{Z})$. We note $G_{2}=G_{1}\star^{r}S\star u_{1}\star\cdots u_{k}$. The edges toggled by an $r$-local complementation over $S$ are described by an element $\omega\in\Omega$. Let us decompose $\omega$ as a linear combination of basis vectors of $\Omega$: $\omega=\omega_{1}+\cdots+\omega_{t}$, where each $\omega_{i}\in\mathcal{B}$. In $G^{\#}_{1}$ and $G^{\#}_{2}$, let $V_{\omega}=\bigcup_{i\in[1,t]}\mathcal{P}^{\omega_{i}}$. Then, $G^{\#}_{2}=G^{\#}_{1}\star^{1}V_{\omega}\star u_{1}\star\cdots u_{k}$. Note that the 1-local complementation over the set $V_{\omega}$ corresponds to the composition of local complementations on each element of $V_{\omega}$. Thus, $G_{1}$ is mapped to $G_{2}$ by a sequence of local complementations that satisfy the additional constraints.
@@
According to Proposition˜3.22, $S$ is genuine. Let $K_{\text{max}}$ be one of the biggest subset (by inclusion) of $V\setminus\textup{{supp}}(S)$ such that $\sum_{N_{G}(u)=K_{\text{max}}}S(u)$ is odd: then $S\bullet\Lambda_{G}^{K_{\text{max}}}$ is odd. If $|V\setminus\textup{{supp}}(S)|\leqslant r+1$, then $|K_{max}|\leqslant r+1$, contradicting the $r$-incidence of $S$. Thus $|V\setminus\textup{{supp}}(S)|\geqslant|K_{max}|\geqslant r+2$.
@@
$=\sum_{N_{G}(u)=K_{0}}S(u)+|\{K\subseteq V\setminus\textup{{supp}}(S)\penalty 10000\ |\penalty 10000\ K_{0}\subsetneq K\}|\bmod 2\text{\penalty 10000\ \penalty 10000\ by hypothesis of induction}$
@@
Let $S^{\prime}$ be the multiset obtained from $S$ by choosing, for every $K\subseteq V\setminus\textup{{supp}}(S)$ s.t. $|K|>1$, a single vertex $u\in\textup{{supp}}(S)$ s.t. $N_{G}(u)=K$, and setting $S^{\prime}(u)=S(u)-1$ (the multiplicity of other vertices remain unchanged). Let us prove that $S^{\prime}$ is $r$-incident and $G\star^{r}S=G\star^{r}S^{\prime}$. First, $S^{\prime}$ is $r$-incident. Indeed, let an integer $k\in[0,r)$, let $K_{1}\subseteq V\setminus\textup{{supp}}(S^{\prime})$ be a set of size $k+2$, and let $k^{\prime}=k-|K_{1}\cap\textup{{supp}}(S)|$. $S\bullet\Lambda_{G}^{K_{1}}$ is a multiple of $2^{r-k^{\prime}-\delta(k^{\prime})}$ by $r$-incidence of $S$, so is $S^{\prime}\bullet\Lambda_{G}^{K_{1}}$, as $S^{\prime}\bullet\Lambda_{G}^{K_{1}}=S\bullet\Lambda_{G}^{K_{1}}-|\{K\subseteq V\setminus\textup{{supp}}(S)\penalty 10000\ |\penalty 10000\ K_{1}\setminus\textup{{supp}}(S)\subseteq K\}|=S\bullet\Lambda_{G}^{K_{1}}-2^{r-k^{\prime}}$. Then, if $u$ or $v$ is in $\textup{{supp}}(S)$, $u\sim_{G\star^{r}S}v\penalty 10000\ \Leftrightarrow\penalty 10000\ u\sim_{G\star^{r}S^{\prime}}v\penalty 10000\ \Leftrightarrow\penalty 10000\ u\sim_{G}v$. If $u,v\in V\setminus\textup{{supp}}(S)$:
$u\sim_{G\star^{r}S}v$ $\penalty 10000\ \Leftrightarrow\penalty 10000\ \left(u\sim_{G}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ S\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
$\penalty 10000\ \Leftrightarrow\penalty 10000\ \left(u\sim_{G}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ S^{\prime}\bullet\Lambda_{G}^{u,v}+2^{r}=2^{r-1}\bmod 2^{r}\right)$
$\penalty 10000\ \Leftrightarrow\penalty 10000\ \left(u\sim_{G}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ S^{\prime}\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)\penalty 10000\ \Leftrightarrow\penalty 10000\ u\sim_{G\star^{r}S^{\prime}}v$
Thus, $G\star^{r}S=G\star^{r}S^{\prime}$. Notice also that $S^{\prime}$ is not genuine. Thus, by Proposition˜3.22 there exists an $S^{\prime\prime}$ such that $G\star^{r-1}S^{\prime\prime}=G\star^{r}S^{\prime}=G\star^{r}S$.